OpenStudy (ledah):

Is the following problem simplified correctly? (5x+7)(5x-7)=25x^2+70x-49 Yes or no?

OpenStudy (ledah):

I Think it is

mathslover (mathslover):

no

Parth (parthkohli):

This is a lot more easier to do than you think hehe. \( \color{Black}{\Rightarrow (a + b) (a - b) = a^2 - b^2}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow (5x)^2 - (7)^2 }\) Can you simplify this?

mathslover (mathslover):

\[\huge{(5x+7)(5x-7) = (5x)^2-7^2 = 25x^2-49}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow (5x)(5x) - (7)(7) }\)

mathslover (mathslover):

just apply the identity that master parth kohli told you

Parth (parthkohli):

Looks like you got the answer right there

OpenStudy (ledah):

So you both told me my answer was wrong that means its not simplified correctly. is that what your saying?